# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ 3.2.1. Case 1
Suppose $v\in\Sigma_{K,\infty}$ or that $v\in\Sigma_{K,f}$ and the $\mathcal{E}_{i,s}$ have good reduction at $v$ (which is to say that the $\mathcal{E}_{i,s,k_{v}}=\mathfrak{G}^{\prime}_{i,\mathfrak{s},k_{v}}$ are elliptic curves). Define the $F_{v}$-vector space
@@
(if $v\in\Sigma_{K,\infty}$) by the canonical isomorphism $R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{s_{0}}\cong R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{s}$ of fibres of the constant sheaf $R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{\Delta_{v}}$.;
(if $v\in\Sigma_{K,f}$) by the pullback $\alpha^{*}_{i}$ associated with an isogeny $\alpha_{i}:\mathcal{E}_{i,s,k_{v}}\to\mathcal{E}_{i,s_{0},k_{v}}$ whose degree we note $d_{i}$.
@@
If $v\in\Sigma_{K,\infty}$, the maps $\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}|_{\Delta_{v}}\to\Delta_{v}$ are smooth and proper, and so the claims follow from [Kat72, Prop. 4.1.2]. For $v\in\Sigma_{K,f}$, the $\sigma_{i}$ are afforded to us by [BO83, (2.4.2)]. Furthermore, in the terminology of [Ogu84, §5], $\mathcal{E}_{i,K_{v}}$ has good reduction over $\mathfrak{C}_{\mathcal{O}_{K_{v}}}$. Therefore, the commutative diagrams in this case are afforded to us by [Ogu84, Rem. 5.14.3] and Lemma 10. ∎
Now we construct relations. To that end, let $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ denote a $K_{v}$-basis for $H^{1}_{DR}(\mathcal{E}_{2,s_{0}}/K_{v})$ whose elements are contained in $\sigma^{-1}_{2}(H^{1}_{v}(\mathcal{E}_{2,s_{0}}))$, and let
@@
Since $f_{0}$ is an isogeny, $\gamma_{v,1,s_{0}}$ and $\delta_{v,2,s_{0}}$ form a $K_{v}$-basis for $H^{1}_{DR}(\mathcal{E}_{1,s_{0}}/K_{v})$.
@@
(and similarly $f_{k_{v}}:\mathcal{E}_{1,s,k_{v}}\to\mathcal{E}_{2,s,k_{v}}$). This proves existence of the diagram. The fact that the diagram commutes in this case is [BO83, (2.4.5)].
The isomorphisms $\epsilon_{i}$ allow us to extend the $\{\gamma_{v,i,s_{0}},\delta_{v,i,s_{0}}\}$ to horizontal bases $\{\gamma_{v,i},\delta_{v,i}\}$ for $H^{1}_{DR}(\mathcal{E}_{i}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ and we let $\Omega_{v,i}\in\mathrm{M}_{2}(\mathcal{O}_{\Delta_{v}})$ be the matrix whose rows give the coordinates of $\{\omega^{v\textrm{-}\mathrm{an}}_{i},\eta^{v\textrm{-}\mathrm{an}}_{i}\}$ in terms of $\{\gamma_{v,i},\delta_{v,i}\}$. By (1) and (2), we have $\Omega_{v,0}:=\Omega_{v,1}(s^{v\textrm{-}\mathrm{an}}_{0})=\Omega_{v,2}(s^{v\textrm{-}\mathrm{an}}_{0})$ and we define $Y_{v,i}:=\Omega_{v,i}\cdot\Omega_{v,0}^{-1}$. By construction, the Taylor series of the entries of the $Y_{v,i}$ with respect to $x$, are precisely the $F_{imn}$ invoked in Section 3.1.
Write
@@
If $v\in\Sigma_{K,\infty}$, the bottom arrow is the extension to $K_{v}$ of an isomorphism $H^{1}_{v}(\mathcal{E}_{2,s})\to H^{1}_{v}(\mathcal{E}_{1,s})$ of $\mathbb{Q}$-vector spaces. By Lemma 12, the sets
@@
For $v\in\Sigma_{K,f}$, we write $f^{\vee}_{0,k_{v}}:\mathcal{E}_{2,s_{0},k_{v}}\to\mathcal{E}_{1,s_{0},k_{v}}$ for the dual isogeny to $f_{0,k_{v}}$. Then, the composition
@@
yields an endomorphism of $\mathcal{E}_{2,s_{0},k_{v}}$ and hence an endomorphism of $H^{1}_{v}(\mathcal{E}_{2,s_{0}})$, which we denote $\phi$. By Lemma 12, this forms part of the commutative diagram
@@
Resuming from the passage preceding Lemma 13, we have
@@
(3) $(F_{2,s})^{-1}A^{T}F_{1,s}=\Omega_{v,0}M^{T}_{v}\Omega_{v,0}^{-1}.$
@@
$\operatorname{tr}((F_{2,s})^{-1}A^{T}F_{1,s})=\operatorname{tr}(M^{T}_{v})=\operatorname{tr}(M_{v}),$
@@ 3.2.2. Case 2
Now suppose $v\in\Sigma_{K,f}$ and that the $\mathcal{E}_{i,s}$ (equivalently, the $\mathcal{E}_{i,s_{0}}$) have bad reduction at $v$. Thus, after possibly replacing $K$ with a finite extension, we have $\mathfrak{G}^{\prime}_{i,\mathfrak{s}_{0},k_{v}}\cong\mathfrak{G}^{\prime}_{i,\mathfrak{s},k_{v}}\cong\mathbb{G}_{m,k_{v}}$. This implies that we can put ourselves in the situation of [DO23, 3.B.1] using the following dictionary, which sends the notations of [DO23, 3.B.1] to those of this section:
$R\mapsto\ \mathcal{O}_{K_{v}},\ \mathfrak{C}\mapsto\ \mathfrak{C}^{\prime}_{\mathcal{O}_{K_{v}}},\ \mathfrak{C}_{0,\mathfrak{p}}\mapsto\ \mathfrak{s}_{0}({\operatorname{Spec}{k_{v}}}),\ \mathfrak{G}\mapsto$ $\ \mathfrak{G}^{\prime}_{\mathcal{O}_{K_{v}}.}$
By [DO23, Lemma 3.3], for some open subspace $\mathcal{C}_{v}$ of $C^{{v\textrm{-}\mathrm{an}}}$ (which was referred to in Section 3.1 (c)) and $\mathcal{E}_{v,i}:=\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}|_{\mathcal{C}_{v}}$, there exist rigid analytic uniformisations $\phi_{v,i}:\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}\times\mathcal{C}_{v}\to\mathcal{E}_{v,i}$. Moreover, by [Ger70, Satz 5], we obtain commutative diagrams
@@
${H^{1}_{DR}(\mathcal{E}_{v,2,s_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,2,s_{0}}}$$\scriptstyle{(f^{v\textrm{-}\mathrm{an}}_{0})^{*}}$${H^{1}_{DR}(\mathcal{E}_{v,1,s_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,1,s_{0}}}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m_{0}]}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})}$    ${H^{1}_{DR}(\mathcal{E}_{v,2,s})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,2,s}}$$\scriptstyle{(f^{v\textrm{-}\mathrm{an}})^{*}}$${H^{1}_{DR}(\mathcal{E}_{v,1,s})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,1,s}}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m]}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}).}$
Consider the natural injective map
$H^{1}_{DR}(\mathcal{E}_{i}/C)^{v\textrm{-}\mathrm{an}}\to H^{1}_{DR}(\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}/C^{v\textrm{-}\mathrm{an}})$
of modules with integrable connection. We also have the pullback
@@
(Indeed, $\mathbb{G}_{m}\times C\to C$ is a rational elementary fibration [ABC20, Def. 25.1.4], hence, its Artin set is equal to $C$ [ABC20, Rmk. 25.2.3]. Therefore, [ABC20, Thm. 32.2.1] gives us this isomorphism.) Since $\Delta_{v}$ is contained in $\mathcal{C}_{v}$, it follows that we can choose a horizontal basis $\{\gamma_{v,2},\delta_{v,2}\}$ for $H^{1}_{DR}(\mathcal{E}_{2}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ such that, under the above morphisms, we have
$\gamma_{v,2}\mapsto dz/z,\ \delta_{v,2}\mapsto 0.$
@@
$\gamma_{v,1,s_{0}}:=(f^{{v\textrm{-}\mathrm{an}}}_{0})^{*}\gamma_{v,1,s_{0}},\ \delta_{v,1,s_{0}}:=(f^{{v\textrm{-}\mathrm{an}}}_{0})^{*}\delta_{v,1,s_{0}}$
for $H^{1}_{DR}(\mathcal{E}_{v,1,s_{0}})$ and extend to a horizontal basis $\{\gamma_{v,1},\delta_{v,1}\}$ for $H^{1}_{DR}(\mathcal{E}_{2}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$. For the above morphisms, we have
$\gamma_{v,1}\mapsto m_{0}\,dz/z,\ \delta_{v,1}\mapsto 0,$
and, with $A$ and $\Omega_{v,i}$ defined as in Section 3.2.1, we arrive at the following equation:
@@
(where we write $(dz/z,0)^{T}$ for the corresponding column vector). If $(\Omega_{v,0})_{21}\neq 0$, we write $\lambda_{v,0}:=(\Omega_{v,0})_{11}(\Omega_{v,0})^{-1}_{21}$ and, expanding, we obtain
(4) $a(F^{v\textrm{-}\mathrm{an}}_{1,s})_{12}+c(F^{v\textrm{-}\mathrm{an}}_{1,s})_{22}-\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{12}$ $=\lambda_{v,0}[\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{11}-a(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-c(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}]$
(5) $b(F^{v\textrm{-}\mathrm{an}}_{1,s})_{12}+d(F^{v\textrm{-}\mathrm{an}}_{1,s})_{22}-\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{22}$ $=\lambda_{v,0}[\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{21}-b(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-d(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}],$
where, again, $(F^{v\textrm{-}\mathrm{an}}_{i,s})_{jk}:=(F^{v\textrm{-}\mathrm{an}}_{i}(x^{v\textrm{-}\mathrm{an}}(s^{v\textrm{-}\mathrm{an}})))_{jk}$. If the right-hand side of (4) is 0, we obtain a linear $v$-adic relation between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. Otherwise, we can divide (5) by (4) and clear denominators to obtain a $v$-adic relation of degree $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. If $(\Omega_{v,0})_{21}=0$, we must have $(\Omega_{v,0})_{11}\neq 0$, and we obtain
(6) $\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{11}-a(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-c(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}=0$
Again, the fact that any of these relations is not in $I$ is readily seen by elementary computation. ∎

### v2

@@ 3.2.1. Case 1
Suppose $v\in\Sigma_{K,\infty}$ or that $v\in\Sigma_{K,f}$ and the $\mathcal{E}_{i,s_{0}}$ have good reduction at $v$ (which is to say that the $\mathcal{E}_{i,s_{0},k_{v}}:=\mathfrak{G}_{i,\mathfrak{s}_{0},k_{v}}$ are elliptic curves).
Setup
Define the $F_{v}$-vector space
@@
(if $v\in\Sigma_{K,\infty}$) by the canonical isomorphism $R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{s_{0}}\cong R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{s}$ of fibres of the constant sheaf $R^{1}(\pi^{v\textrm{-}\mathrm{an}}_{i})_{*}\mathbb{Q}|_{\Delta_{v}}$;
(if $v\in\Sigma_{K,f}$) by the pullback $\alpha^{*}_{i}$ associated with an isogeny $\alpha_{i}:\mathcal{E}_{i,s,k_{v}}\to\mathcal{E}_{i,s_{0},k_{v}}$ whose degree we denote $d_{i}$.
@@
If $v\in\Sigma_{K,\infty}$, the maps $\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}|_{\Delta_{v}}\to\Delta_{v}$ are smooth and proper, and so the claims follow from [Kat72, Prop. 4.1.2].
If $v\in\Sigma_{K,f}$, the $\sigma_{i}$ are afforded to us by [BO83, Cor. 2.5]. Furthermore, in the terminology of [Ogu84, §5], $\mathcal{E}_{i,K_{v}}$ has good reduction over $\mathfrak{C}_{\mathcal{O}_{K_{v}}}$. Since $\Delta_{v}$ is simple, the $\mathcal{O}_{K}$-extensions of $s$ and $s_{0}$ have the same reduction at $v$. Therefore, the commutative diagrams in this case are afforded to us by [Ogu84, Rem. 5.14.3]. ∎
Let $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ denote a $K_{v}$-basis for $H^{1}_{DR}(\mathcal{E}_{2,s_{0}}/K_{v})$ whose elements are contained in the $F_{v}$-subspace $\sigma^{-1}_{2}(H^{1}_{v}(\mathcal{E}_{2,s_{0}}))$, and let
@@
Since $f_{0}$ is a $K$-isogeny, $\gamma_{v,1,s_{0}}$ and $\delta_{v,2,s_{0}}$ form a $K_{v}$-basis for $H^{1}_{DR}(\mathcal{E}_{1,s_{0}}/K_{v})$.
@@
(and similarly $f_{k_{v}}:\mathcal{E}_{1,s,k_{v}}\to\mathcal{E}_{2,s,k_{v}}$). This induces the morphism along the bottom of the diagram. The fact that the diagram commutes in this case is [BO83, (2.4.5)].
The parallel transport isomorphisms $\epsilon_{i}$ allow us to extend the $\{\gamma_{v,i,s_{0}},\delta_{v,i,s_{0}}\}$ to horizontal bases $\{\gamma_{v,i},\delta_{v,i}\}$ for $H^{1}_{DR}(\mathcal{E}_{i}/C_{0})^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ and we let $\Omega_{v,i}\in\mathrm{M}_{2}(\mathcal{O}_{\Delta_{v}})$ be the matrix whose rows give the coordinates of $\{\omega^{v\textrm{-}\mathrm{an}}_{i},\eta^{v\textrm{-}\mathrm{an}}_{i}\}$ in terms of $\{\gamma_{v,i},\delta_{v,i}\}$. We define $\theta_{v,i}:=\Omega_{v,i}(s^{v\textrm{-}\mathrm{an}}_{0})$ and $Y_{v,i}:=\Omega_{v,i}\cdot\theta_{v,i}^{-1}$. By construction, the Taylor series of the entries of the $Y_{v,i}$ with respect to $x$, are precisely the $F_{imn}$ invoked in Section 3.1. By (3) and (4), we have
(5) $\theta_{v,1}=\mu\theta_{v,2}.$
Remark 3.4.
The argument above establishes that the $F_{imn}$ have $v$-adic radius of convergence at least $1$ at almost all $v\in\Sigma_{K}$. By Lemma 2.1, they have positive $v$-adic radius of convergence at the finitely many remaining $v\in\Sigma_{K}$. This may look similar to the proof of [And89, Thm. B] in [And89, V, App.], but this is not sufficient to establish that the $F_{imn}$ are G-functions. The proof in [And89, V, App.] uses the stronger fact that the $v$-adic generic radius of convergence of the differential equation satisfied by the $F_{imn}$ is $1$ for almost all $v\in\Sigma_{K}$.
Relations
The set-up above did not use the existence of an isogeny $f\colon\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}$. Now we use this isogeny to construct relations between evaluations of elements of $\mathcal{G}$.
To that end, write
@@
If $v\in\Sigma_{K,\infty}$, the bottom arrow is the extension to $K_{v}$ of an isomorphism $H^{1}_{v}(\mathcal{E}_{2,s})\to H^{1}_{v}(\mathcal{E}_{1,s})$ of $\mathbb{Q}$-vector spaces. By Lemma 3.3, the sets
@@
For $v\in\Sigma_{K,f}$, we write $f^{\vee}_{0,k_{v}}:\mathcal{E}_{2,s_{0},k_{v}}\to\mathcal{E}_{1,s_{0},k_{v}}$ for the dual isogeny to $f_{0,k_{v}}$. Then, using the isogenies $\alpha_{1}$, $\alpha_{2}$ from Lemma 3.3, the composition
@@
yields an endomorphism of $\mathcal{E}_{2,s_{0},k_{v}}$ and hence an endomorphism of $H^{1}_{v}(\mathcal{E}_{2,s_{0}})$, which we denote $\phi$. By Lemma 3.3 and the naturality of the $\sigma_{i}$, we obtain the commutative diagram
@@
Resuming from the passage preceding Lemma 3.5, we have
@@
$A^{T}F_{1,s}\theta_{v,1}=F_{2,s}\theta_{v,2}M_{v}^{T}.$
Using (5), this rearranges to
(6) $\mu(F_{2,s})^{-1}A^{T}F_{1,s}=\theta_{v,2}M^{T}_{v}\theta_{v,2}^{-1}.$
@@
$\operatorname{tr}(\mu(F_{2,s})^{-1}A^{T}F_{1,s})=\operatorname{tr}(M^{T}_{v})=\operatorname{tr}(M_{v}),$
@@ 3.2.2. Case 2
Now suppose $v\in\Sigma_{K,f}$ and that the $\mathcal{E}_{i,s_{0}}$ (equivalently, the $\mathcal{E}_{i,s}$) have bad reduction at $v$. Thus, after possibly replacing $K$ with a finite extension, we have $\mathfrak{G}_{i,\mathfrak{s}_{0},k_{v}}\cong\mathfrak{G}_{i,\mathfrak{s},k_{v}}\cong\mathbb{G}_{m,k_{v}}$. At such a place, we may interpret the $v$-adic evaluations of the G-functions associated with $\mathcal{E}_{i}\to C_{0}$ in a similar way to the “locally invariant periods” around a point of multiplicative degeneration, as in [DO25] or [DO23].
More precisely, we can put ourselves in the situation of [DO23, 3.B.1] using the following dictionary, which sends the notations of [DO23, 3.B.1] to those of this section:
(7) $R\mapsto\ \mathcal{O}_{K_{v}},\ \mathfrak{C}\mapsto\ \mathfrak{C}_{\mathcal{O}_{K_{v}}},\ \mathfrak{C}_{0,\mathfrak{p}}\mapsto\ \mathfrak{s}_{0}(\mathfrak{p}_{v}),\ \mathfrak{G}\mapsto$ $\ \mathfrak{G}_{\mathcal{O}_{K_{v}}.}$
Let $\mathcal{C}_{v}$ denote the residue disc $\operatorname{red}_{v}^{-1}(\mathfrak{s}_{0}(\mathfrak{p}_{v}))\subset\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}^{\mathrm{rig}}\subset C^{v\textrm{-}\mathrm{an}}$ (where, as before, $\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}$ denotes the formal completion of $\mathfrak{C}$ along $\mathfrak{C}_{k_{v}}$). By [Ber96, Prop. 0.2.7], $\mathcal{C}_{v}$ is equal to the image $\hat{\mathfrak{C}}_{\mathfrak{s}_{0}(\mathfrak{p}_{v})}^{\mathrm{rig}}\subset C^{v\textrm{-}\mathrm{an}}$ of the formal completion $\hat{\mathfrak{C}}_{\mathfrak{s}_{0}(\mathfrak{p}_{v})}$ of $\mathfrak{C}$ along $\mathfrak{s}_{0}(\mathfrak{p}_{v})$ under Berthelot’s rigid generic fibre functor. Thus, via the dictionary (7), the rigid space denoted $\mathcal{C}=\mathfrak{C}_{\mathrm{for}{\dagger}}^{\mathrm{rig}}$ in [DO23, 3.B.1] corresponds to $\mathcal{C}_{v}$.
Let $\mathcal{E}_{v,i}:=\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}|_{\mathcal{C}_{v}}$. By [DO23, Prop. 3.1 and Lemma 3.3], there exist rigid analytic uniformisations $\phi_{v,i}:\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}\times\mathcal{C}_{v}\to\mathcal{E}_{v,i}$. (Note that we do not require any compatibility between the rigid uniformisations at different places $v$, essentially because the functions $Y_{v}$ defined in Lemma 2.3 are independent of the choice of horizontal sections $\gamma_{v,i}$.) Moreover, by [Ger70, Satz 5], we obtain commutative diagrams
@@
(8) ${H^{1}_{DR}(\mathcal{E}_{v,2,s_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,2,s_{0}}}$$\scriptstyle{(f^{v\textrm{-}\mathrm{an}}_{0})^{*}}$${H^{1}_{DR}(\mathcal{E}_{v,1,s_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,1,s_{0}}}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m_{0}]}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})}$    ${H^{1}_{DR}(\mathcal{E}_{v,2,s})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,2,s}}$$\scriptstyle{(f^{v\textrm{-}\mathrm{an}})^{*}}$${H^{1}_{DR}(\mathcal{E}_{v,1,s})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}_{v,1,s}}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m]}$${H^{1}_{DR}(\mathbb{G}^{v\textrm{-}\mathrm{an}}_{m}).}$
Since $\mathcal{E}_{i}\to C_{0}$ is proper, the natural map
$H^{1}_{DR}(\mathcal{E}_{i}/C_{0})^{v\textrm{-}\mathrm{an}}\to H^{1}_{DR}(\mathcal{E}^{v\textrm{-}\mathrm{an}}_{i}/C^{v\textrm{-}\mathrm{an}}_{0})$
of modules with integrable connection is an isomorphism by standard applications of rigid analytic GAGA (or, more precisely, [Con06b, Appendix A, (A.1.1)]). We also have the pullback
@@
(Indeed, $\mathbb{G}_{m}\times C_{0}\to C_{0}$ is a rational elementary fibration [ABC20, Def. 25.1.4], hence, its Artin set is equal to $C_{0}$ [ABC20, Rmk. 25.2.3]. Therefore, [ABC20, Thm. 32.2.1] gives us this isomorphism.)
Since $\Delta_{v}$ is simple, we have $\Delta_{v}\subset\mathcal{C}_{v}$. By Lemma 2.1, after shrinking $\Delta_{v}$ and $R_{v}$, we may assume that $H^{1}_{DR}(\mathcal{E}_{2}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ possesses a basis of horizontal sections. It follows that the restriction of $\phi_{v,2}^{*}|_{\Delta_{v}}$ to horizontal sections is a non-zero linear map from a $K_{v}$-vector space of dimension $2$ to a $K_{v}$-vector space of dimension $1$. Therefore, we can choose a horizontal basis $\{\gamma_{v,2},\delta_{v,2}\}$ for $H^{1}_{DR}(\mathcal{E}_{2}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ such that
$\phi_{v,2}^{*}(\gamma_{v,2})=dz/z,\quad\phi_{v,2}^{*}(\delta_{v,2})=0.$
@@
$\gamma_{v,1,s_{0}}:=(f^{{v\textrm{-}\mathrm{an}}}_{0})^{*}\gamma_{v,2,s_{0}},\ \delta_{v,1,s_{0}}:=(f^{{v\textrm{-}\mathrm{an}}}_{0})^{*}\delta_{v,2,s_{0}}$
for $H^{1}_{DR}(\mathcal{E}_{v,1,s_{0}})$. Again by Lemma 3.3, after shrinking $\Delta_{v}$ and $R_{v}$, we may assume that $H^{1}_{DR}(\mathcal{E}_{1}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ possesses a basis of horizontal sections, so that $\{\gamma_{v,1,s_{0}},\delta_{v,1,s_{0}}\}$ extends to a horizontal basis $\{\gamma_{v,1},\delta_{v,1}\}$ for $H^{1}_{DR}(\mathcal{E}_{1}/C_{0})^{v\textrm{-}\mathrm{an}}(\Delta_{v})$. By (8) and the fact that horizontal sections form a local system, we have
$\phi_{v,1}^{*}(\gamma_{v,1})=m_{0}\,dz/z,\quad\phi_{v,1}^{*}(\delta_{v,1})=0.$
Using the right-hand diagram from (8), with $A$ and $\Omega_{v,i}$ defined as in Section 3.2.1, we arrive at the following equation:
@@
(where we write $(dz/z,0)^{T}$ for the corresponding column vector). If $(\theta_{v,2})_{21}\neq 0$, we write $\lambda_{v,2}:=(\theta_{v,2})_{11}(\theta_{v,2})^{-1}_{21}$ and, expanding, we obtain
(9) $a\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{12}+c\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{22}-\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{12}$ $=\lambda_{v,2}[\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{11}-a\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-c\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}]$
(10) $b\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{12}+d\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{22}-\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{22}$ $=\lambda_{v,2}[\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{21}-b\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-d\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}],$
where, again, $(F^{v\textrm{-}\mathrm{an}}_{i,s})_{jk}:=(F^{v\textrm{-}\mathrm{an}}_{i}(x^{v\textrm{-}\mathrm{an}}(s^{v\textrm{-}\mathrm{an}})))_{jk}$. If the right-hand side of (9) is 0, we obtain a linear $v$-adic relation between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. Otherwise, we can divide (10) by (9) and clear denominators to obtain a $v$-adic relation of degree $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$. If $(\theta_{v,2})_{21}=0$, we must have $(\theta_{v,2})_{11}\neq 0$, and we obtain
(11) $\frac{m}{m_{0}}(F^{v\textrm{-}\mathrm{an}}_{2,s})_{11}-a\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{11}-c\mu(F^{v\textrm{-}\mathrm{an}}_{1,s})_{21}=0$
Again, the fact that these relations are not in $I$ is readily seen by elementary computation. ∎